I need to solve this problem
log5x/logx= 3
could you please help me solve this equation
thank you very much, viki

Answer by Alan3354(6097)

log5x/logx= 3
--------------
log(5x) = 3log(x)
log(5x) = log(x^3)
x^3 = 5x
x^3 - 5x = 0
x(x^2-5) = 0
x = 0
0 is not allowed
-----------------
x = +sqrt(5)
------------
-sqrt(5) is not allowed due to the log function